A chemist is trying to make 30 gallons of a 42 percent solution. He only has a 50 percent and a 60 percent solution. How much of each are needed?
Answer by Alan3354(69443)

A chemist is trying to make 30 gallons of a 42 percent solution. He only has a 50 percent and a 60 percent solution. How much of each are needed?
---------------
You can only find 10000 examples of this problem on this site.
---------
f = amount of 50%
s = amount of 60%
----
f + s = 30
50f + 60s = 42*(f + s)